OpenStudy (anonymous):

I have been having a hard time with this please Help. Graph each pair of parametric equations 1..x = 9in t + sin 8t

OpenStudy (phi):

with a graphing calculator or by hand ?

OpenStudy (anonymous):

any i guess

OpenStudy (anonymous):

i jus need to kno how to solve it

OpenStudy (phi):

t will be an angle ranging from 0 to 2pi (or 0º to 360º) you would pick angles in that interval, figure out x and y, plot the point, and continue. When you have enough "dots", connect the dots to make a curve.

OpenStudy (phi):

It is a tedious process by hand

OpenStudy (phi):

For example, when t=0 you would compute x = 7 sin t + sin 7t x = 7 sin(0) + sin(7*0) = 0 y = 7 cos t + cos 7t y = 7 cos(0) + cos(7*0) = 7*1+1= 8 plot the point(0,8) now try other (close by) angles, and plot the points

OpenStudy (lastdaywork):

I was trying to make a differential equation; this is how far I got - 7x = 42sint - dy/dt 7y = 42cost + dx/dt I'll try after dinner.

OpenStudy (anonymous):

thank you!!